# GiGL: Large-Scale Graph Neural Networks at Snapchat

### v1

@@ Abstract.
Recent advances in graph machine learning (ML) with the introduction of Graph Neural Networks (GNNs) have led to a widespread interest in applying these approaches to business applications at scale. GNNs enable differentiable end-to-end (E2E) learning of model parameters given graph structure which enables optimization towards popular node, edge (link) and graph-level tasks. While the research innovation in new GNN layers and training strategies has been rapid, industrial adoption and utility of GNNs has lagged considerably due to the unique scale challenges that large-scale graph ML problems create. In this work, we share our approach to training, inference, and utilization of GNNs at Snapchat. To this end, we present GiGL (Gigantic Graph Learning), an open-source library to enable large-scale distributed graph ML to the benefit of researchers, ML engineers, and practitioners. We use GiGL internally at Snapchat to manage the heavy lifting of GNN workflows, including graph data preprocessing from relational DBs, subgraph sampling, distributed training, inference, and orchestration. GiGL is designed to interface cleanly with open-source GNN modeling libraries prominent in academia like PyTorch Geometric (PyG), while handling scaling and productionization challenges that make it easier for internal practitioners to focus on modeling. GiGL is used in multiple production settings, and has powered over 35 launches across multiple business domains in the last 2 years in the contexts of friend recommendation, content recommendation and advertising. This work details high-level design and tools the library provides, scaling properties, case studies in diverse business settings with large-scale graphs up to hundreds of millions of nodes, tens of billions of edges, and hundreds of node and edge features, and several key lessons learned in employing graph ML at scale on large social data. GiGL is open-sourced at https://github.com/snap-research/GiGL.
@@ 1. Introduction
We built GiGL to address our internal needs over the years. GiGL currently powers all GNN-based graph machine learning workflows at Snapchat, supporting critical product features such as friend recommendations, lens recommendations, spam and abuse detection, and content recommendations. Across multiple use-cases, we use GiGL to generate either task-specific embeddings or general-purpose user embeddings, which serve as inputs for embedding-based dense retrieval (EBR) systems and/or as added dense-features in lower-funnel ranking models. Yet, we believe there are opportunities for GiGL to serve a broader community. GiGL addresses a critical need for machine learning researchers, engineers, and practitioners: enabling large-scale exploration and experimentation with state-of-the-art GNN models on industry-scale graphs, while compatibly interfacing with familiar open-source libraries which enable fast model iteration and development like PyG. While these libraries are more commonly used in contexts with few millions of nodes and edges, they benefit from strong community support and cutting-edge GNN research implementations. GiGL bridges this gap by providing tooling around handling distributed graph scaling challenges, utilities, orchestration and more to handle heavy lifting, allowing researchers to continue leveraging the modeling tools they know and trust. Hence, this work accompanies an open-source release of GiGL, at https://github.com/snap-research/GiGL. We will evolve this offering to support the broader graph ML community, and are happy to accept contributions.
@@ 2. Preliminaries and Related Work
Most GNN applications in the research landscape occur at a single-machine scale, owing to the size of benchmark datasets (Hu et al., 2020). In industrial contexts, graphs may surpass such limits in terms of number of nodes, edges and features, which complicates GNN training and inference due to data dependency in the message-passing design and the need for careful minibatching (Duan et al., 2022; Xue et al., 2024a). Recent literature discusses multiple directions to train GNNs with large-scale graphs, including sampling methods (Hamilton et al., 2017; Chen et al., 2018), staleness-based methods (Fey et al., 2021; Xue et al., 2024b), distillation (Zhang et al., 2021; Guo et al., 2023), quantization (Ding et al., 2021), condensation and coarsening (Jin et al., 2021; Tsitsulin and Perozzi, 2023), pre-training (Han et al., 2022b; Borisyuk et al., 2024), distributed training (Md et al., 2021; Lin et al., 2023), and many more (Xue et al., 2024a). These techniques have been commonly adopted as directions in GNN scaling research, and are applied to training in larger benchmark graphs (e.g., ogbn-products (Hu et al., 2020) with 2.4M nodes and 61M edges). Nonetheless, in large-scale applications, some of these techniques meet various challenges owing to inefficient implementations, misalignment between academic and industrial assumptions, and diverse graph properties (Palowitch et al., 2022) – for example, graphs we handle at Snapchat are routinely have 10-50B edges.
@@ 3. Design and Interface
GiGL is designed with horizontal scaling across many compute resources in mind, making it well-suited for deployment on both custom compute clusters and cloud environments. While vertically-scaled GNN solutions (hosting graph and training on one very large machine) are viable to an extent, their scalability is inherently limited111We discuss our experience moving away from a vertically-scaled approach in Section 6.1., and is a bottleneck in environments where data volume can fluctuate across tasks, and grow rapidly with user growth.
@@ 3.1. Components and Orchestration
We begin by introducing the tabularization scheme, which has been a core part of GiGL and has been in use in multiple domains inside Snapchat with significant success over the past two years. Figure 2 provides a visual overview of GiGL’s workflow for a sample link prediction task, trained with a triplet-style contrastive loss. We next outline the design of GiGL pipelines, which are comprised of five main components and orchestration capabilities222Detailed documentation of each component can be found at https://snap-research.github.io/GiGL/#gigl-components:
@@
(DP) is a distributed feature transformation pipeline backed by TensorFlow Transform (TFT)333https://www.tensorflow.org/tfx/transform, leveraging the Beam444https://beam.apache.org runner. TFT is widely used in industry for large-scale data preprocessing, encompassing feature scaling, normalization, handling of categorical features, and more. DP reads and transforms input graph topology (edge-lists) and raw node and edge features from relational data sources (in our case, BigQuery or Google Cloud Storage) into the TFRecord format. Users are able to define distributed transforms like one-hot-encoding, normalization, scaling, imputation and filtering over the graph. We provide bindings to run DP on Google Cloud Dataflow555https://cloud.google.com/products/dataflow (with AWS support planned).
@@
A common practitioner workflow is the configuration of an end-to-end pipeline which runs components in sequence (see Figure 2) to power batch inference. GiGL leverages workflow management software to schedule the component jobs in order. GiGL currently provides tooling to launch pipelines on Kubeflow666https://www.kubeflow.org and VertexAI777https://cloud.google.com/vertex-ai. For easier orchestration, we introduce two additional mini-components which run before pipeline-start: (i) a Validation Checker which verifies input configuration correctness and path-existence to avoid wasting development and compute cycles, and (ii) a Config Populator, which takes template configurations, and populates it with frozen asset paths (transformed data, splits, model artifact, inferences) to simplify and centralize asset tracking. Path persistence prior to pipeline execution enables retrying individual components idempotently, fault tolerance, and amortization of individual component outputs for other runs: e.g., multiple Trainers using the same SG output during hyperparameter tuning.
@@ 3.2. Real-time Subgraph Sampling
While tabularization provides a scalable and effective approach for subgraph generation, it has some limits (discussed in Section 6.1) when greater flexibility and adaptive use of a graph is required during training. Real-time subgraph sampling workflows enable this flexibility for speed, at the expense of amortization potential. To enable such workflows, we integrate and build from a custom version of GraphLearn-for-PyTorch888https://github.com/alibaba/graphlearn-for-pytorch (GLT). While sharing the same orchestration, data reading and transformation (DP) logic, our customized GLT backend partitions the graph into a distributed in-memory dataset across machines. The graphs are semi-randomly partitioned: nodes are shuffled across machines, and adjacent edges are collocated based on (customizable) adjacent source or destination node. Hence, all nodes are able to access all 1-hop neighbors within-machine, and the associated node and edge features are also collocated. Once this dataset is established, training and inference workflows are able to generate subgraphs via gRPC calls across machines for multi-hop sampling. During training, we adopt a similar strategy as in SG to enable globally consistent distributed splits to mask out relevant data, and can flexibly use CPU or GPU training. This logic operates within the logical scope of the Trainer and Inferencer components, and similar to the tabularization scheme: trainer outputs a model artifact and inference outputs final embeddings or predictions to a sink (BigQuery, or Google Cloud Storage, internally). We have successfully used this scheme internally, and open-source support is currently work-in-progress.
@@ 3.3. User Interface
Apart from the raw relational data, users are asked to specify two configuration files which dictate the GiGL workflow which runs: a Resource Config and a Task Config; see Appendix C for links to examples of them, respectively, and our official user guide999https://snap-research.github.io/GiGL/#configuration for further details and guidelines.
@@ 4. Experimentation and Scaling
Table 1 showcases the component-level runtime comparison of GiGL running on the internal graph and one public benchmark, MAG240M101010We also offer example codes for running MAG240M end-to-end with GiGL at: https://github.com/snap-research/GiGL/tree/main/examples/MAG240M (Hu et al., 2021). We can observe that GiGL is able to finish end-to-end on the full industry-scale graph within 12 hours, making it feasible for product use-cases. Additionally, from the time comparison of the different versions, we can observe that the number of edges generally has a bigger impact on the runtime, aligning with GNNs’ time complexity which is linear with the number of edges (Wu et al., 2020). Beyond these graph scaling experiments, we also conducted additional offline experiments of GiGL with varying resources and setups, which are included in Tables 8 and 9 in Appendix B.
@@ 5.2. Content Recommendation
Beyond friend recommendation, content recommendation is another key application with abundant relational data that benefits from GNN modeling. The interactions between users and short videos (Spotlight111111https://www.snapchat.com/spotlight and Discover121212https://www.snapchat.com/discover tab) naturally form bipartite graphs, with user nodes connected to video nodes through various interactions. By incorporating additional metadata, such as content creators, this structure extends to a heterogeneous graph, effectively representing relationship across multiple types of entities.
@@ 6.1. Lessons: Infrastructure
Our early explorations in GNN modeling at industry-scale occurred when modern GNN libraries were limited, and “spatial GNNs” were still emerging (Hamilton et al., 2017; Ying et al., 2018). To handle large graphs, we employed a vertical-scaling solution, leveraging the highest-memory cloud machines available at the time (GCP’s n1-highmem-96, with 96 vCPU and 624GB RAM). We stored graph topology in CSR format, node features as a large tensor, and used Ray 131313https://github.com/ray-project/ray for both shared-memory and multiprocessing, enabling a producer-consumer setup across CPU workers and GPU trainers.
@@
In our early efforts, libraries like PyG (Fey and Lenssen, 2019) and DGL (Wang et al., 2019) were nascent, and not well-suited for large-scale industrial use. Hence, we developed graph sampling and convolution operations in native PyTorch, using models like GraphSAGE (Hamilton et al., 2017) and (neighbor-sampled) GCN (Kipf and Welling, 2016). This approach required managing complex multi-dimensional tensors141414E.g., tensor with shape of (batch_size, n_nodes, n_nodes, n_node_feat) for 2-layer convolution. and intricate reshaping for message passing. While effective initially, these custom solutions were challenging to adapt to advanced GNN models. Hence, when re-architecting, we take advantage of advances in frameworks like PyG and DGL which abstract away low-level convolution details. GiGL now benefits from the continuous improvements contributed to these libraries, closing the gap between cutting-edge research and production applications.
@@ 6.2. Lessons: Modeling
As with many ML applications, offline-online discrepancies are common. While we use metrics like validation loss, NDCG, and HitRate to guide model development, we find that improving offline metrics has diminishing returns for online performance after a point, due to external factors. Hence, we often employ post-processing logic to also monitor task-relevant (non-differentiable) surrogate metrics which can help prune unpromising offline candidates early.
@@ 6.3. Opportunities
Standard message-passing neural networks (MPNNs) are not the final frontier for link prediction. More complex subgraph-GNN architectures (Zhang and Chen, 2018; Zhu et al., 2021b; Zhao et al., 2022a) show superior link performance, and newer work show strong effectiveness of joint usage of simple heuristic link prediction metrics with GNNs (Mao et al., 2024; Ma et al., 2024). Moreover, co-training of large language models with graphs to improve node embedding quality is an emerging area (Chen et al., 2024; Ren et al., 2024). We see explorations in all these areas as avenues for the next-generation of GNNs at Snapchat.
@@ Acknowledgements.
- Ching et al. (2015) Avery Ching, Sergey Edunov, Maja Kabiljo, Dionysios Logothetis, and Sambavi Muthukrishnan. 2015. One trillion edges: Graph processing at facebook-scale. Proceedings of the VLDB Endowment 8, 12 (2015), 1804–1815.
- Derrow-Pinion et al. (2021) Austin Derrow-Pinion, Jennifer She, David Wong, Oliver Lange, Todd Hester, Luis Perez, Marc Nunkesser, Seongjae Lee, Xueying Guo, Brett Wiltshire, et al. 2021. Eta prediction with graph neural networks in google maps. In Proceedings of the 30th ACM international conference on information & knowledge management. 3767–3776.
@@
- Fan et al. (2019) Wenqi Fan, Yao Ma, Qing Li, Yuan He, Eric Zhao, Jiliang Tang, and Dawei Yin. 2019. Graph neural networks for social recommendation. In The world wide web conference. 417–426.
@@
- Fey et al. (2021) Matthias Fey, Jan E Lenssen, Frank Weichert, and Jure Leskovec. 2021. Gnnautoscale: Scalable and expressive graph neural networks via historical embeddings. In International conference on machine learning. PMLR, 3294–3304.
@@
- Guo et al. (2023) Zhichun Guo, William Shiao, Shichang Zhang, Yozen Liu, Nitesh V Chawla, Neil Shah, and Tong Zhao. 2023. Linkless link prediction via relational distillation. In International Conference on Machine Learning. PMLR, 12012–12033.
@@
- Hamilton et al. (2017) Will Hamilton, Zhitao Ying, and Jure Leskovec. 2017. Inductive representation learning on large graphs. Advances in neural information processing systems 30 (2017).
@@
- Hu et al. (2020) Weihua Hu, Matthias Fey, Marinka Zitnik, Yuxiao Dong, Hongyu Ren, Bowen Liu, Michele Catasta, and Jure Leskovec. 2020. Open graph benchmark: Datasets for machine learning on graphs. Advances in neural information processing systems 33 (2020), 22118–22133.
- Jain et al. (2019) Ankit Jain, Isaac Liu, Ankur Sarda, and Piero Molino. 2019. Food discovery with uber eats: Using graph learning to power recommendations. Accessed March 1 (2019), 2021.
@@
- Jin et al. (2020) Wei Jin, Yao Ma, Xiaorui Liu, Xianfeng Tang, Suhang Wang, and Jiliang Tang. 2020. Graph structure learning for robust graph neural networks. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 66–74.
@@
- Ma et al. (2024) Li Ma, Haoyu Han, Juanhui Li, Harry Shomer, Hui Liu, Xiaofeng Gao, and Jiliang Tang. 2024. Mixture of Link Predictors on Graphs. In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
- Ma et al. (2021) Yao Ma, Xiaorui Liu, Tong Zhao, Yozen Liu, Jiliang Tang, and Neil Shah. 2021. A unified view on graph neural networks as graph signal denoising. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management. 1202–1211.
@@
- Shi et al. (2023) Jiahui Shi, Vivek Chaurasiya, Yozen Liu, Shubham Vij, Yan Wu, Satya Kanduri, Neil Shah, Peicheng Yu, Nik Srivastava, Lei Shi, et al. 2023. Embedding Based Retrieval in Friend Recommendation. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval. 3330–3334.
@@
- Tang et al. (2022) Xianfeng Tang, Yozen Liu, Xinran He, Suhang Wang, and Neil Shah. 2022. Friend story ranking with edge-contextual local graph convolutions. In Proceedings of the Fifteenth ACM International Conference on Web Search and Data Mining. 1007–1015.
- Tang et al. (2020) Xianfeng Tang, Yozen Liu, Neil Shah, Xiaolin Shi, Prasenjit Mitra, and Suhang Wang. 2020. Knowing your fate: Friendship, action and temporal explanations for user engagement prediction on social apps. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 2269–2279.
@@
- Wang et al. (2017) Ruoxi Wang, Bin Fu, Gang Fu, and Mingliang Wang. 2017. Deep & cross network for ad click predictions. Proceedings of the ADKDD’17 (2017), 1–7.
@@
- Wu et al. (2024) Jiancan Wu, Xiang Wang, Xingyu Gao, Jiawei Chen, Hongcheng Fu, and Tianyu Qiu. 2024. On the effectiveness of sampled softmax loss for item recommendation. ACM Transactions on Information Systems 42, 4 (2024), 1–26.
@@
- Yang et al. (2020) Carl Yang, Aditya Pal, Andrew Zhai, Nikil Pancha, Jiawei Han, Charles Rosenberg, and Jure Leskovec. 2020. MultiSage: Empowering GCN with contextualized multi-embeddings on web-scale multipartite networks. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 2434–2443.
@@
- Zhao et al. (2022b) Tong Zhao, Wei Jin, Yozen Liu, Yingheng Wang, Gang Liu, Stephan Günnemann, Neil Shah, and Meng Jiang. 2022b. Graph Data Augmentation for Graph Machine Learning: A Survey. Data Engineering (2022), 140.
- Zhao et al. (2021) Tong Zhao, Yozen Liu, Leonardo Neves, Oliver Woodford, Meng Jiang, and Neil Shah. 2021. Data augmentation for graph neural networks. In Proceedings of the aaai conference on artificial intelligence, Vol. 35. 11015–11023.

### v2

@@ Abstract.
Recent advances in graph machine learning (ML) with the introduction of Graph Neural Networks (GNNs) have led to a widespread interest in applying these approaches to business applications at scale. GNNs enable differentiable end-to-end (E2E) learning of model parameters given graph structure which enables optimization towards popular node, edge (link) and graph-level tasks. While the research innovation in new GNN layers and training strategies has been rapid, industrial adoption and utility of GNNs has lagged considerably due to the unique scale challenges that large-scale graph ML problems create. In this work, we share our approach to training, inference, and utilization of GNNs at Snapchat. To this end, we present GiGL (Gigantic Graph Learning), an open-source library to enable large-scale distributed graph ML to the benefit of researchers, ML engineers, and practitioners. We use GiGL internally at Snapchat to manage the heavy lifting of GNN workflows, including graph data preprocessing from relational DBs, subgraph sampling, distributed training, inference, and orchestration. GiGL is designed to interface cleanly with open-source GNN modeling libraries prominent in academia like PyTorch Geometric (PyG), while handling scaling and productionization challenges that make it easier for internal practitioners to focus on modeling. GiGL is used in multiple production settings, and has powered over 35 launches across multiple business domains in the last 2 years in the contexts of friend recommendation, content recommendation and advertising. This work details high-level design and tools the library provides, scaling properties, case studies in diverse business settings with large-scale graphs up to hundreds of millions of nodes, tens of billions of edges, and hundreds of node and edge features, and several key lessons learned in employing graph ML at scale on large social data. GiGL is open-sourced at https://github.com/Snapchat/GiGL.
@@ 1. Introduction
We built GiGL to address our internal needs over the years. GiGL currently powers all GNN-based graph machine learning workflows at Snapchat, supporting critical product features such as friend recommendations, lens recommendations, spam and abuse detection, and content recommendations. Across multiple use-cases, we use GiGL to generate either task-specific embeddings or general-purpose user embeddings, which serve as inputs for embedding-based dense retrieval (EBR) systems and/or as added dense-features in lower-funnel ranking models. Yet, we believe there are opportunities for GiGL to serve a broader community. GiGL addresses a critical need for machine learning researchers, engineers, and practitioners: enabling large-scale exploration and experimentation with state-of-the-art GNN models on industry-scale graphs, while compatibly interfacing with familiar open-source libraries which enable fast model iteration and development like PyG. While these libraries are more commonly used in contexts with few millions of nodes and edges, they benefit from strong community support and cutting-edge GNN research implementations. GiGL bridges this gap by providing tooling around handling distributed graph scaling challenges, utilities, orchestration and more to handle heavy lifting, allowing researchers to continue leveraging the modeling tools they know and trust. Hence, this work accompanies an open-source release of GiGL111https://github.com/Snapchat/GiGL. We will evolve this offering to support the broader graph ML community, and are happy to accept contributions.
@@ 2. Preliminaries and Related Work
Most GNN applications in the research landscape occur at a single-machine scale, owing to the size of benchmark datasets (Hu et al., 2020a). In industrial contexts, graphs may surpass such limits in terms of number of nodes, edges and features, which complicates GNN training and inference due to data dependency in the message-passing design and the need for careful minibatching (Duan et al., 2022; Xue et al., 2024a). Recent literature discusses multiple directions to train GNNs with large-scale graphs, including sampling methods (Hamilton et al., 2017; Chen et al., 2018), staleness-based methods (Fey et al., 2021; Xue et al., 2024b), distillation (Zhang et al., 2021; Guo et al., 2023), quantization (Ding et al., 2021), condensation and coarsening (Jin et al., 2021; Tsitsulin and Perozzi, 2023), pre-training (Han et al., 2022b; Borisyuk et al., 2024), distributed training (Md et al., 2021; Lin et al., 2023), and many more (Xue et al., 2024a). These techniques have been commonly adopted as directions in GNN scaling research, and are applied to training in larger benchmark graphs (e.g., ogbn-products (Hu et al., 2020a) with 2.4M nodes and 61M edges). Nonetheless, in large-scale applications, some of these techniques meet various challenges owing to inefficient implementations, misalignment between academic and industrial assumptions, and diverse graph properties (Palowitch et al., 2022) – for example, graphs we handle at Snapchat are routinely have 10-50B edges.
@@ 3. Design and Interface
GiGL is designed with horizontal scaling across many compute resources in mind, making it well-suited for deployment on both custom compute clusters and cloud environments. While vertically-scaled GNN solutions (hosting graph and training on one very large machine) are viable to an extent, their scalability is inherently limited222We discuss our experience moving away from a vertically-scaled approach in Section 6.1., and is a bottleneck in environments where data volume can fluctuate across tasks, and grow rapidly with user growth.
@@ 3.1. Components and Orchestration
We begin by introducing the tabularization scheme, which has been a core part of GiGL and has been in use in multiple domains inside Snapchat with significant success over the past two years. Figure 2 provides a visual overview of GiGL’s workflow for a sample link prediction task, trained with a triplet-style contrastive loss. We next outline the design of GiGL pipelines, which are comprised of five main components and orchestration capabilities333Detailed documentation of each component can be found at https://snapchat.github.io/GiGL/docs/user_guide/overview/architecture.html:
@@
(DP) is a distributed feature transformation pipeline backed by TensorFlow Transform (TFT)444https://www.tensorflow.org/tfx/transform, leveraging the Beam555https://beam.apache.org runner. TFT is widely used in industry for large-scale data preprocessing, encompassing feature scaling, normalization, handling of categorical features, and more. DP reads and transforms input graph topology (edge-lists) and raw node and edge features from relational data sources (in our case, BigQuery or Google Cloud Storage) into the TFRecord format. Users are able to define distributed transforms like one-hot-encoding, normalization, scaling, imputation and filtering over the graph. We provide bindings to run DP on Google Cloud Dataflow666https://cloud.google.com/products/dataflow (with AWS support planned).
@@
A common practitioner workflow is the configuration of an end-to-end pipeline which runs components in sequence (see Figure 2) to power batch inference. GiGL leverages workflow management software to schedule the component jobs in order. GiGL currently provides tooling to launch pipelines on Kubeflow777https://www.kubeflow.org and VertexAI888https://cloud.google.com/vertex-ai. For easier orchestration, we introduce two additional mini-components which run before pipeline-start: (i) a Validation Checker which verifies input configuration correctness and path-existence to avoid wasting development and compute cycles, and (ii) a Config Populator, which takes template configurations, and populates it with frozen asset paths (transformed data, splits, model artifact, inferences) to simplify and centralize asset tracking. Path persistence prior to pipeline execution enables retrying individual components idempotently, fault tolerance, and amortization of individual component outputs for other runs: e.g., multiple Trainers using the same SG output during hyperparameter tuning.
@@ 3.2. Real-time Subgraph Sampling
While tabularization provides a scalable and effective approach for subgraph generation, it has some limits (discussed in Section 6.1) when greater flexibility and adaptive use of a graph is required during training. Real-time subgraph sampling workflows enable this flexibility for speed, at the expense of amortization potential. To enable such workflows, we integrate and build from a custom version of GraphLearn-for-PyTorch999https://github.com/alibaba/graphlearn-for-pytorch (GLT). While sharing the same orchestration, data reading and transformation (DP) logic, our customized GLT backend partitions the graph into a distributed in-memory dataset across machines. The graphs are semi-randomly partitioned: nodes are shuffled across machines, and adjacent edges are collocated based on (customizable) adjacent source or destination node. Hence, all nodes are able to access all 1-hop neighbors within-machine, and the associated node and edge features are also collocated. Once this dataset is established, training and inference workflows are able to generate subgraphs via gRPC calls across machines for multi-hop sampling. During training, we adopt a similar strategy as in SG to enable globally consistent distributed splits to mask out relevant data, and can flexibly use CPU or GPU training. This logic operates within the logical scope of the Trainer and Inferencer components, and similar to the tabularization scheme: trainer outputs a model artifact and inference outputs final embeddings or predictions to a sink (BigQuery, or Google Cloud Storage, internally). We have successfully used this scheme internally, and open-source support is currently work-in-progress.
@@ 3.3. User Interface
Apart from the raw relational data, users are asked to specify two configuration files which dictate the GiGL workflow which runs: a Resource Config and a Task Config; see Appendix C for links to examples of them, respectively, and our official user guide101010https://snapchat.github.io/GiGL/#configuration for further details and guidelines.
@@ 4. Experimentation and Scaling
Table 1 showcases the component-level runtime comparison of GiGL running on the internal graph and one public benchmark, MAG240M111111We also offer example codes for running MAG240M end-to-end with GiGL at: https://github.com/Snapchat/GiGL/tree/main/examples/MAG240M (Hu et al., 2021). We can observe that GiGL is able to finish end-to-end on the full industry-scale graph within 12 hours, making it feasible for product use-cases. Additionally, from the time comparison of the different versions, we can observe that the number of edges generally has a bigger impact on the runtime, aligning with GNNs’ time complexity which is linear with the number of edges (Wu et al., 2020). Beyond these graph scaling experiments, we also conducted additional offline experiments of GiGL with varying resources and setups, which are included in Tables 9 and 7 in Appendix B.
@@ 5.2. Content Recommendation
Beyond friend recommendation, content recommendation is another key application with abundant relational data that benefits from GNN modeling. The interactions between users and short videos (Spotlight121212https://www.snapchat.com/spotlight and Discover131313https://www.snapchat.com/discover tab) naturally form bipartite graphs, with user nodes connected to video nodes through various interactions. By incorporating additional metadata, such as content creators, this structure extends to a heterogeneous graph, effectively representing relationship across multiple types of entities.
@@ 6.1. Lessons: Infrastructure
Our early explorations in GNN modeling at industry-scale occurred when modern GNN libraries were limited, and “spatial GNNs” were still emerging (Hamilton et al., 2017; Ying et al., 2018). To handle large graphs, we employed a vertical-scaling solution, leveraging the highest-memory cloud machines available at the time (GCP’s n1-highmem-96, with 96 vCPU and 624GB RAM). We stored graph topology in CSR format, node features as a large tensor, and used Ray 141414https://github.com/ray-project/ray for both shared-memory and multiprocessing, enabling a producer-consumer setup across CPU workers and GPU trainers.
@@
In our early efforts, libraries like PyG (Fey and Lenssen, 2019) and DGL (Wang et al., 2019) were nascent, and not well-suited for large-scale industrial use. Hence, we developed graph sampling and convolution operations in native PyTorch, using models like GraphSAGE (Hamilton et al., 2017) and (neighbor-sampled) GCN (Kipf and Welling, 2016). This approach required managing complex multi-dimensional tensors151515E.g., tensor with shape of (batch_size, n_nodes, n_nodes, n_node_feat) for 2-layer convolution. and intricate reshaping for message passing. While effective initially, these custom solutions were challenging to adapt to advanced GNN models. Hence, when re-architecting, we take advantage of advances in frameworks like PyG and DGL which abstract away low-level convolution details. GiGL now benefits from the continuous improvements contributed to these libraries, closing the gap between cutting-edge research and production applications.
@@ 6.2. Lessons: Modeling
As with many ML applications, offline-online discrepancies are common. Hence, we often employ post-processing logic to also monitor task-relevant (non-differentiable) surrogate metrics which can help prune unpromising offline candidates early.
@@ 6.3. Opportunities
Standard message-passing neural networks (MPNNs) are not the final frontier for link prediction. More complex subgraph-GNN architectures (Zhang and Chen, 2018; Zhu et al., 2021b; Zhao et al., 2022a) show superior link performance, and newer work show strong effectiveness of joint usage of simple heuristic link prediction metrics with GNNs (Mao et al., 2024; Ma et al., 2024; Zhao et al., 2022c). Moreover, co-training of large language models with graphs to improve node embedding quality is an emerging area (Chen et al., 2024; Ren et al., 2024; Zhao et al., 2024). We see explorations in all these areas as avenues for the next-generation of GNNs at Snapchat.
@@ Acknowledgements.
- Ching et al. (2015) Avery Ching, Sergey Edunov, Maja Kabiljo, Dionysios Logothetis, and Sambavi Muthukrishnan. 2015. One trillion edges: Graph processing at facebook-scale. Proceedings of the VLDB Endowment (2015).
- Derrow-Pinion et al. (2021) Austin Derrow-Pinion, Jennifer She, David Wong, Oliver Lange, Todd Hester, Luis Perez, Marc Nunkesser, Seongjae Lee, Xueying Guo, Brett Wiltshire, et al. 2021. Eta prediction with graph neural networks in google maps. In Proceedings of the 30th ACM international conference on information & knowledge management.
@@
- Fan et al. (2019) Wenqi Fan, Yao Ma, Qing Li, Yuan He, Eric Zhao, Jiliang Tang, and Dawei Yin. 2019. Graph neural networks for social recommendation. In The world wide web conference.
@@
- Fey et al. (2021) Matthias Fey, Jan E Lenssen, Frank Weichert, and Jure Leskovec. 2021. Gnnautoscale: Scalable and expressive graph neural networks via historical embeddings. In International conference on machine learning.
@@
- Guo et al. (2023) Zhichun Guo, William Shiao, Shichang Zhang, Yozen Liu, Nitesh V Chawla, Neil Shah, and Tong Zhao. 2023. Linkless link prediction via relational distillation. In International Conference on Machine Learning.
@@
- Hamilton et al. (2017) Will Hamilton, Zhitao Ying, and Jure Leskovec. 2017. Inductive representation learning on large graphs. In Advances in neural information processing systems.
@@
- Hu et al. (2020a) Weihua Hu, Matthias Fey, Marinka Zitnik, Yuxiao Dong, Hongyu Ren, Bowen Liu, Michele Catasta, and Jure Leskovec. 2020a. Open graph benchmark: Datasets for machine learning on graphs. Advances in neural information processing systems 33 (2020), 22118–22133.
- Hu et al. (2020b) Weihua Hu, Matthias Fey, Marinka Zitnik, Yuxiao Dong, Hongyu Ren, Bowen Liu, Michele Catasta, and Jure Leskovec. 2020b. Open Graph Benchmark: Datasets for Machine Learning on Graphs. arXiv preprint arXiv:2005.00687 (2020).
- Jain et al. (2019) Ankit Jain, Isaac Liu, Ankur Sarda, and Piero Molino. 2019. Food discovery with uber eats: Using graph learning to power recommendations. Accessed (2019).
@@
- Jin et al. (2020) Wei Jin, Yao Ma, Xiaorui Liu, Xianfeng Tang, Suhang Wang, and Jiliang Tang. 2020. Graph structure learning for robust graph neural networks. In ACM SIGKDD international conference on knowledge discovery & data mining.
@@
- Ma et al. (2024) Li Ma, Haoyu Han, Juanhui Li, Harry Shomer, Hui Liu, Xiaofeng Gao, and Jiliang Tang. 2024. Mixture of Link Predictors on Graphs. In Annual Conference on Neural Information Processing Systems.
- Ma et al. (2021) Yao Ma, Xiaorui Liu, Tong Zhao, Yozen Liu, Jiliang Tang, and Neil Shah. 2021. A unified view on graph neural networks as graph signal denoising. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management.
@@
- Shi et al. (2023) Jiahui Shi, Vivek Chaurasiya, Yozen Liu, Shubham Vij, Yan Wu, Satya Kanduri, Neil Shah, Peicheng Yu, Nik Srivastava, Lei Shi, et al. 2023. Embedding Based Retrieval in Friend Recommendation. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval.
@@
- Tang et al. (2022) Xianfeng Tang, Yozen Liu, Xinran He, Suhang Wang, and Neil Shah. 2022. Friend story ranking with edge-contextual local graph convolutions. In Proceedings of the Fifteenth ACM International Conference on Web Search and Data Mining.
- Tang et al. (2020) Xianfeng Tang, Yozen Liu, Neil Shah, Xiaolin Shi, Prasenjit Mitra, and Suhang Wang. 2020. Knowing your fate: Friendship, action and temporal explanations for user engagement prediction on social apps. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining.
@@
- Wang et al. (2017) Ruoxi Wang, Bin Fu, Gang Fu, and Mingliang Wang. 2017. Deep & cross network for ad click predictions. Proceedings of the ADKDD’17 (2017).
@@
- Wu et al. (2024) Jiancan Wu, Xiang Wang, Xingyu Gao, Jiawei Chen, Hongcheng Fu, and Tianyu Qiu. 2024. On the effectiveness of sampled softmax loss for item recommendation. ACM Transactions on Information Systems (2024).
@@
- Yang et al. (2020) Carl Yang, Aditya Pal, Andrew Zhai, Nikil Pancha, Jiawei Han, Charles Rosenberg, and Jure Leskovec. 2020. MultiSage: Empowering GCN with contextualized multi-embeddings on web-scale multipartite networks. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining.
@@
- Zhao et al. (2022b) Tong Zhao, Wei Jin, Yozen Liu, Yingheng Wang, Gang Liu, Stephan Günnemann, Neil Shah, and Meng Jiang. 2022b. Graph Data Augmentation for Graph Machine Learning: A Survey. Data Engineering (2022).
- Zhao et al. (2022c) Tong Zhao, Gang Liu, Daheng Wang, Wenhao Yu, and Meng Jiang. 2022c. Learning from counterfactual links for link prediction. In International Conference on Machine Learning.
- Zhao et al. (2021) Tong Zhao, Yozen Liu, Leonardo Neves, Oliver Woodford, Meng Jiang, and Neil Shah. 2021. Data augmentation for graph neural networks. In Proceedings of the aaai conference on artificial intelligence.
- Zhao et al. (2024) Tong Zhao, Neil Shah, and Elham Ghazizadeh. 2024. Learning from graphs beyond message passing neural networks. In The Second Tiny Papers Track at ICLR 2024.